# Entanglement and apparent thermality in simulated black holes

### v1

@@ 3 Quantum field theory of spinor in Rindler space-time
To quantise the solution of Eq. (C.31) to the Dirac equation for a spinor in Rindler space-time, $\psi(\tilde{t},\rho)$ and $\pi(\tilde{t},\rho)$ are promoted to Hermitian operators $\hat{\psi}(\tilde{t},\rho)$ and $\hat{\pi}(\tilde{t},\rho)$ and required to satisfy the equal-time canonical anti-commutation relations of Eq. (B.19). Using these anti-commutation relations and the canonical momentum $\hat{\pi}(\tilde{t},\rho)=-\frac{i}{2}\hat{\psi}^{\dagger}(\tilde{t},\rho)$ conjugate to $\hat{\psi}(\tilde{t},\rho)$, which is given in Eq. (B.18) for a general Riemann-Cartan space-time, the following anti-commutation relation between the field operator $\hat{\psi}(\tilde{t},\rho)$ and its Hermitian conjugate $\hat{\psi}^{\dagger}(\tilde{t},\rho)$ can be obtained
$\{\hat{\psi}_{\alpha}(\tilde{t},\rho),\hat{\psi}^{\dagger}_{\beta}(\tilde{t},\rho^{\prime})\}=\delta_{\alpha\beta}\delta(\rho-\rho^{\prime}),$ (C.33)
where $\alpha,\beta$ denote spinor components. In quantising the Dirac field, the functions $\tilde{f}(k)$ and $\tilde{g}(k)$ are also promoted to operators $\hat{f}(k)$ and $\hat{g}(k)$. Using the inverse of the eigenfunction decomposition defined in Eq. (C.28), it can be verified that these operators can be expressed in terms of the Cauchy initial data, $\hat{\psi}(0,\rho)$ and $\hat{\pi}(0,\rho)$, as $\hat{f}(k)=\langle\hat{\psi}_{k}(\rho),\hat{\psi}(0,\rho)\rangle$ and $\hat{g}(k)=\langle\hat{\phi}_{k}(\rho),\hat{\psi}(0,\rho)\rangle$ [57]. Using these expressions for the operators $\hat{f}(k)$ and $\hat{g}(k)$ and the orthonormality conditions of $\psi_{k}(\rho)$ and $\phi_{k}(\rho)$, the following anti-commutation relations can be derived: [57]
$\begin{split}\{\hat{f}(k),\hat{f}^{\dagger}(q)\}=\{&\hat{g}(k),\hat{g}^{\dagger}(q)\}=\delta(k-q),\\ \{\hat{f}(k),\hat{f}(q)\}&=\{\hat{g}(k),\hat{g}(q)\}=0,\end{split}$ (C.34)
It can be explicitly shown that $\hat{f}(k)$ and $\hat{g}^{\dagger}(k)$ are creation and annihilation operators that act to raise and lower the energy of the Dirac field by discrete values. To do this, the Hamiltonian of the Dirac spinor on the Rindler space-time must be expressed in terms of $\hat{f}(k)$ and $\hat{g}^{\dagger}(k)$. From (B.22) for the Hamiltonian of a spinor on a stationary Riemann-Cartan space-time, the Hamiltonian of the spinor on the Rindler space-time will be
$\hat{H}=\int d\rho\ \hat{\pi}(\tilde{t},\rho)\partial_{t}\hat{\psi}(\tilde{t},\rho),$ (C.35)
@@
$\hat{H}=-\frac{i}{2}\int d\rho(\partial_{t}\hat{\psi}^{\dagger}(\tilde{t},\rho))\hat{\psi}(\tilde{t},\rho).$ (C.36)
Substitution of Eq. (C.31) for the general solution to the Dirac equation in a Rindler space-time into the above Hamiltonian, and making use of the orthonormality conditions of $\psi_{k}(\rho)$ and $\phi_{k}(\rho)$, yields [57]
$\hat{H}=\frac{|k|}{2}\int d\mu(k)\left(\hat{f}^{\dagger}(k)\hat{f}(k)-\hat{g}^{\dagger}(k)\hat{g}(k)\right).$ (C.37)
From the Hamiltonian of a Dirac spinor in a Rindler space-time given in Eq. (C.37) and the anti-commutation relations of the operators $\hat{f}(k)$ and $\hat{g}(k)$ given in Eq. (C.34), the following commutation relations between $\hat{f}(k)$, $\hat{g}(k)$ and the Hamiltonian can be obtained:
$[\hat{H},\hat{f}^{\dagger}(k)]=\frac{|k|}{2}\hat{f}^{\dagger}(k),\ \ \ [\hat{H},\hat{f}(k)]=-\frac{|k|}{2}\hat{f}(k),$ (C.38)
$[\hat{H},\hat{g}^{\dagger}(k)]=-\frac{|k|}{2}\hat{g}^{\dagger}(k),\ \ \ [\hat{H},\hat{g}(k)]=\frac{|k|}{2}\hat{g}(k),$ (C.39)
which, in conjunction with the eigenvalue equations of $\psi_{k}(k)$ and $\phi_{k}(k)$ given in Eq. (C.32), can be used to deduce that $\hat{f}^{\dagger}(k)$ and $\hat{f}(k)$ are operators that act to raise and lower the energy of the quantum field by creating and annihilating quanta $\psi_{k}$ of energy $|k|/2$, respectively, whilst $\hat{g}^{\dagger}(k)$ and $\hat{g}(k)$ are operators that act to lower and raise the energy of the quantum field by creating and annihilating quanta $\phi_{k}$ of energy |k|/2-|k|/2, respectively. To be in line with standard notation, we re-label $\hat{f}^{\dagger}(k),\hat{f}(k)\mapsto\hat{a}^{\dagger}_{k},\hat{a}_{k}$, and $\hat{g}^{\dagger}(k),\hat{g}(k)\mapsto\hat{b}_{k},\hat{b}^{\dagger}_{k}$. Therefore, the general solution to the Dirac equation in Rindler space-time is [51, 49]
$\hat{\psi}(\tilde{t},\rho)=\int^{\infty}_{-\infty}d\mu(k)\left(\hat{a}_{k}e^{-i|k|\tilde{t}}\psi_{k}(\rho)+\hat{b}^{\dagger}_{k}e^{i|k|\tilde{t}}\phi_{k}(\rho)\right),$ (C.40)
where $\hat{a}_{k}=\langle\psi_{k}(\rho),\psi(0,\rho)\rangle$, $\hat{b}^{\dagger}_{k}=\langle\phi_{k}(\rho),\psi(0,\rho)\rangle$, and $\{\hat{a}_{k},\hat{a}^{\dagger}_{q}\}=\{\hat{b}_{k},\hat{b}^{\dagger}_{q}\}=\delta(k-q)$, whilst all other anti-commutation relations are zero.
@@ 4 Quantum field theory of spinor in Minkowski space-time
In Minkowski space-time, as the zweibein is trivially given by $e^{a}_{\mu}=\mathbb{I}$, the Dirac equation, as is given in Eq. (C.15), simplifies to
$\left(i\gamma^{a}e^{\mu}_{a}\partial_{\mu}-m\right)\Psi(T,X)=0$ (C.43)
for the two-component spinor $\Psi(T,X)$, which, after expanding the Einstein summation over $a$ and $\mu$, reduces to
$\left(i\gamma^{0}\partial_{T}+i\gamma^{1}\partial_{X}-m\right)\Psi(T,X)=0.$ (C.44)
@@
Analogous to the Rindler metric, as the Minkowski metric is stationary with respect to the Minkowski time coordinate $T$, the vector $K^{\mu}=\delta^{\mu}_{T}$ (or equivalently $\partial_{T}$) is a Killing vector that is time-like for all of the Minkowski space-time. Thus, the energy $\tilde{E}$ of the spinor is a well-defined quantity that is conserved along the Minkowski metric’s geodesics [57, 23], and the Dirac equation can be solved via the separation of variables
$\Psi(T,X)=e^{-i\tilde{E}_{K}T}\Psi_{\tilde{E}_{K}}(X),$ (C.45)
@@
where $\Psi_{\tilde{E}_{K}}(X)=(\Psi_{\tilde{E}_{K}}^{+}(X)\ \Psi_{\tilde{E}_{K}}^{-}(X))^{T}$ is a two-component spinor that depends only on $X$. Therefore, the Minkowski space-time Dirac equation can be expressed as the following eigenvalue equation:
$\tilde{E}_{K}\gamma^{0}\Psi_{\tilde{E}_{K}}(X)+i\gamma^{1}\partial_{X}\Psi_{\tilde{E}_{K}}(X)-m\Psi_{\tilde{E}_{K}}(K)=0,$ (C.46)
@@
where $h=i\gamma^{0}\gamma^{1}\partial_{X}+m\gamma^{0}$ is a linear differential operator that is Hermitian under the conserved inner product
$\langle\Phi(X),\Psi(X)\rangle=\int d\mu(X)\Phi^{\dagger}(X)\Psi(X),$ (C.47)
@@
and $d\mu(X)=dX$ is a measure of integration. That is, $h$ satisfies $\langle\Phi,h\Psi\rangle=\langle h\Phi,\Psi\rangle$ [65]. Once again restricting ourselves to the case of a massless spinor and choosing the chiral representation of the $\gamma$ matrices, the Hermitian operator $h$ can be expressed as $h=-u\sigma^{z}\partial_{X}\Psi_{\tilde{E}_{K}}(X)$. Therefore, the eigenvalue equation corresponds to the two following decoupled differential equations:
$\tilde{E}_{K}\Psi_{\tilde{E}_{K}}^{\pm}(X)\pm i\partial_{X}\Psi_{\tilde{E}_{K}}^{\pm}(x)=0,$ (C.48)
@@
where $\Psi_{\tilde{E}_{K}}^{\pm}(x)$ denote the components of the spinor $\Psi_{\tilde{E}_{K}}(x)$ that have the solutions $\Psi_{\tilde{E}_{K}}^{\pm}(X)=\mathcal{N}e^{\pm i\tilde{E}_{K}X}$, where $\mathcal{N}=1/\sqrt{2\pi}$ is a normalisation constant chosen such that $\langle\Psi_{\tilde{E}_{K}}(X),\Psi_{\tilde{E}_{Q}}(X)\rangle=\delta(K-Q)$. Hence, as a solution to the Dirac equation in Minkowski space-time, we have
$\Psi(T,X)=\frac{1}{\sqrt{2\pi}}e^{-i\tilde{E}_{K}T}\begin{pmatrix}e^{i\tilde{E}_{K}X}\\ e^{-i\tilde{E}_{K}X}\end{pmatrix}$ (C.49)
@@
The charge conjugate solution $\Phi(T,X)=C\Psi^{*}(T,X)=e^{i\tilde{E}T}\Phi_{\tilde{E}_{K}}(X)$ to the Dirac equation in Minkowski space-time, which satisfies the eigenvalue equation $h\Phi_{\tilde{E}_{K}}(x)=-\tilde{E}_{K}\Phi_{\tilde{E}_{K}}(X)$ and the orthonormality condition $\langle\Phi_{\tilde{E}_{K}},\Psi_{K}\rangle=0$, is given by
$\Phi(T,X)=\Psi^{*}(T,X)=\frac{1}{\sqrt{2\pi}}e^{i\tilde{E}_{K}T}\begin{pmatrix}e^{-iKT}\\ e^{iKT}\end{pmatrix}$. (C.50)
@@
Again, assuming that the orthonormal basis formed by the eigenfunctions $\Psi_{\tilde{E}_{K}}(X)$ of the Hamiltonian $h$ is complete, any function $F(X)$ within the Hilbert space $\mathcal{H}_{M}$ defined by the conserved inner product of Eq. (C.47) can be decomposed into a continuous sum of eigenfunctions $\Psi_{\tilde{E}_{K}}(X)$ as [53, 65, 64]
$F(X)=\int d\mu(K)\tilde{f}(K)\Psi_{\tilde{E}_{K}}(X),$ (C.51)
@@
$\langle\tilde{f}(K),\tilde{g}(K)\rangle=\int\mu(K)\tilde{f}^{\dagger}(K)\tilde{g}(K),$ (C.52)
@@
of a corresponding Hilbert space $\mathcal{L}^{2}(K)$ of functions $\tilde{f}(K)$, which is related to $\mathcal{H}_{M}$ via the eigenfunction decomposition of Eq. (C.51) and its inverse [53, 65, 64]
$\tilde{f}(K)=\int d\mu(X)\Psi_{\tilde{E}_{K}}^{\dagger}(X)F(X).$ (C.53)
@@
For a consistent correspondence between the two Hilbert spaces via the eigenfunction decomposition and its inverse, we require the normalisation conditions $\langle\Psi_{\tilde{E}_{K}}(X),\Psi_{Q}(X)\rangle=\delta(K-Q)$, where $\int d\mu(Q)\delta(K-Q)\tilde{f}(Q)=\tilde{f}(K)$, and $\langle\Psi_{\tilde{E}_{K}}(X),\Psi_{\tilde{E}_{K}}(X^{\prime})\rangle=\delta(X-X^{\prime})$, where $\int d\mu(X^{\prime})\delta(X-X^{\prime})F(X^{\prime})=F(X)$ [53, 65, 64].
@@
Using the eigenfunction decomposition defined in Eq. (C.51), the most general solution $\Psi(T,X)$ of the Minkowski space-time Dirac equation can be decomposed into the continuous sum of both the spinor $\Psi_{\tilde{E}_{K}}(X)$ and charge conjugate spinor $\Phi_{\tilde{E}_{K}}(X)$ eigenfunctions of the Dirac equation. That is,
$\begin{split}\Psi(T,X)=\int^{\infty}_{-\infty}d\mu(K)\Bigl(&\tilde{f}(K)e^{-i\tilde{E}_{K}T}\Psi_{\tilde{E}_{K}}(X)\\ &+\tilde{g}(K)e^{i\tilde{E}_{K}T}\Phi_{\tilde{E}_{K}}(X)\Bigr).\end{split}$ (C.54)
@@
Once again, to ensure that the spinor $\Psi_{\tilde{E}_{K}}$ and its charge conjugate $\Phi_{\tilde{E}_{K}}$ always correspond to positive and negative energy solutions, respectively, we define
$K=\begin{cases}\tilde{E}_{K},&\text{if }\tilde{E}_{K}\geq 0\\ -\tilde{E}_{K},&\text{if }\tilde{E}_{K}<0\end{cases},\ u_{K}=\begin{cases}u_{+},&\text{if }K\geq 0\\ u_{-},&\text{if}K<0\end{cases}.$ (C.55)
such that the general solution to the Dirac equation in Eq. (C.56) can be expressed as
$\begin{split}\Psi(T,X)=\int^{\infty}_{-\infty}d\mu(K)\Bigl(&\tilde{f}(K)e^{-i|K|T}\Psi_{K}(X)\\ &+\tilde{g}(K)e^{i|K|T}\Phi_{K}(X)\Bigr),\end{split}$ (C.56)
where $\Psi_{K}(X)=\mathcal{N}e^{-iKX}u_{K}$ and $\Phi_{K}(X)=\mathcal{N}e^{iKX}u_{K}$.
@@
Once again, this solution to the Dirac equation for a spinor in Minkowski space-time can be quantised by promoting $\Psi(T,X)$ and $\Pi(T,X)=-\frac{i}{2}\Psi^{\dagger}(T,X)$ to Hermitian operators and imposing the equal-time canonical anti-commutations given in Eq. (B.19). In quantising the Dirac field in Minkowski space-time, the functions $\tilde{f}(K)$ and $\tilde{g}(K)$ are promoted to the operators $\hat{f}(K)$ and $\hat{g}(K)$, respectively. Using the inverse eigenfunction decomposition defined in (C.53), it can be verified that the operators $\hat{f}(K)$ and $\hat{g}(K)$ can be expressed in terms of the Cauchy initial data, $\hat{\Psi}(0,X)$ and $\hat{\Pi}(0,X)$, as $\hat{f}(K)=\langle\hat{\Psi}_{K}(X),\hat{\Psi}(0,X)\rangle$ and $\hat{g}(K)=\langle\hat{\Phi}_{K}(X),\hat{\Psi}(0,X)\rangle$. Using these expressions for the operators $\hat{f}(K)$ and $\hat{g}(K)$ and the orthonormality conditions of $\hat{\Psi}_{K}(X)$ and $\hat{\Phi}_{K}(X)$, the following anti-commutation relations can be derived [57]:
$\{\hat{f}(K),\hat{f}^{\dagger}(Q)\}=\{\hat{g}(K),\hat{g}^{\dagger}(Q)\}=\delta(K-Q),$ (C.57)
$\{\hat{f}(K),\hat{f}(Q)\}=\{\hat{g}(K),\hat{g}^{\dagger}(Q)\}=0,$ (C.58)
@@
Analogous to the Rindler space-time case, it then follows that $\hat{f}^{\dagger}(K)$ and $\hat{g}(K)$ are operators that act to raise and lower the energy of the quantum field by creating and annihilating quanta $\Psi_{K}$ of energy $|K|/2$, respectively, whilst $\hat{g}^{\dagger}(K)$ and $\hat{g}(K)$ are operators that act to lower and raise the energy of the quantum field by creating and annihilating quanta $\Phi_{K}$ of energy |K|/2-|K|/2, respectively. To be in line with standard notation, we re-label $\hat{f}(K),\hat{f}(K)\mapsto\hat{A}^{\dagger}(K),\hat{A}_{K}$, and $\hat{g}^{\dagger}(K),\hat{g}(K)\mapsto\hat{B}_{K},\hat{B}^{\dagger}_{K}$. Therefore, the general solution to the Dirac equation in Minkowski space-time is [49]
$\begin{split}\hat{\Psi}(T,X)=\int^{\infty}_{-\infty}d\mu(K)\Bigl(&\hat{A}_{K}e^{-i|K|T}\Psi_{K}(X)\\ &+\hat{B}^{\dagger}_{K}e^{i|K|T}\Phi_{K}(X)\Bigr),\end{split}$ (C.59)
where $\hat{A}_{K}=\langle\Psi_{K},\Psi(0,X)\rangle$, $\hat{B}^{\dagger}_{K}=\langle\Phi_{K},\Psi(0,X)\rangle$, and $\{\hat{A}_{K},\hat{A}^{\dagger}_{Q}\}=\{\hat{B}_{K},\hat{B}^{\dagger}_{Q}\}=\delta(K-Q)$, whilst all other anti-commutation relations are zero.

### v2

@@ 3 Quantum field theory of spinor in Rindler space-time
To quantise the solution of Eq. (B.28) to the Dirac equation in Rindler space-time, $\psi(\tilde{t},\rho)$ and $\pi(\tilde{t},\rho)$ are promoted to Hermitian operators $\hat{\psi}(\tilde{t},\rho)$ and $\hat{\pi}(\tilde{t},\rho)$ and required to satisfy the equal-time canonical anti-commutation relations of Eq. (A.19), where the canonical momentum conjugate to $\hat{\psi}(\tilde{t},\rho)$ is given by $\hat{\pi}(\tilde{t},\rho)=-\frac{i}{2}\hat{\psi}^{\dagger}(\tilde{t},\rho)$ for the Rindler metric and by Eq. (A.18) for a general Riemann-Cartan space-time. In quantising the Dirac field, the functions $\tilde{f}(k)$ and $\tilde{g}(k)$ are also promoted to operators $\hat{f}(k)$ and $\hat{g}(k)$. Using the inverse of the eigenfunction decomposition defined in Eq. (B.25), it can be verified that these operators can be expressed in terms of the Cauchy initial data, $\hat{\psi}(0,\rho)$ and $\hat{\pi}(0,\rho)$, as $\hat{f}(k)=\langle\hat{\psi}_{k}(\rho),\hat{\psi}(0,\rho)\rangle$ and $\hat{g}(k)=\langle\hat{\phi}_{k}(\rho),\hat{\psi}(0,\rho)\rangle$ [48]. Using these expressions and the orthonormality conditions of $\psi_{k}(\rho)$ and $\phi_{k}(\rho)$, the following anti-commutation relations can be derived: [48]
$\begin{split}\{\hat{f}(k),\hat{f}^{\dagger}(q)\}=\{&\hat{g}(k),\hat{g}^{\dagger}(q)\}=\delta(k-q),\\ \{\hat{f}(k),\hat{f}(q)\}&=\{\hat{g}(k),\hat{g}(q)\}=0,\end{split}$ (B.30)
which are those satisfied by typical creation and annihilation operators.
It can be explicitly shown that $\hat{f}(k)$ and $\hat{g}^{\dagger}(k)$ are creation and annihilation operators that act to raise and lower the energy of the Dirac field by discrete values. To do this, the Hamiltonian of the Dirac spinor on the Rindler space-time must be expressed in terms of $\hat{f}(k)$ and $\hat{g}^{\dagger}(k)$. From (A.22) for the Hamiltonian of a spinor on a stationary Riemann-Cartan space-time, the Hamiltonian of the spinor on the Rindler space-time will be
$\hat{H}=\int d\rho\ \hat{\pi}(\tilde{t},\rho)\partial_{t}\hat{\psi}(\tilde{t},\rho),$ (B.31)
@@
$\hat{H}=-\frac{i}{2}\int d\rho(\partial_{t}\hat{\psi}^{\dagger}(\tilde{t},\rho))\hat{\psi}(\tilde{t},\rho).$ (B.32)
Substitution of Eq. (B.28) for the general solution to the Dirac equation in a Rindler space-time into the above Hamiltonian, and making use of the orthonormality conditions of $\psi_{k}(\rho)$ and $\phi_{k}(\rho)$, yields [48]
$\hat{H}=\frac{|k|}{2}\int d\mu(k)\left(\hat{f}^{\dagger}(k)\hat{f}(k)-\hat{g}^{\dagger}(k)\hat{g}(k)\right).$ (B.33)
From this Hamiltonian and the anti-commutation relations of the operators $\hat{f}(k)$ and $\hat{g}(k)$ given in Eq. (B.30), the following commutation relations between $\hat{f}(k)$, $\hat{g}(k)$ and the Hamiltonian can be obtained:
$[\hat{H},\hat{f}^{\dagger}(k)]=\frac{|k|}{2}\hat{f}^{\dagger}(k),\ \ \ [\hat{H},\hat{f}(k)]=-\frac{|k|}{2}\hat{f}(k),$ (B.34)
$[\hat{H},\hat{g}^{\dagger}(k)]=-\frac{|k|}{2}\hat{g}^{\dagger}(k),\ \ \ [\hat{H},\hat{g}(k)]=\frac{|k|}{2}\hat{g}(k),$ (B.35)
which, in conjunction with the eigenvalue equations of $\psi_{k}(k)$ and $\phi_{k}(k)$ given in Eq. (B.29), can be used to deduce that $\hat{f}^{\dagger}(k)$ and $\hat{f}(k)$ are operators that act to raise and lower the energy of the quantum field by creating and annihilating quanta $\psi_{k}$ of energy $|k|/2$, respectively, whilst $\hat{g}^{\dagger}(k)$ and $\hat{g}(k)$ act to lower and raise the energy of the quantum field by creating and annihilating quanta $\phi_{k}$ of energy |k|/2-|k|/2, respectively. To be in line with standard notation, we re-label $\hat{f}^{\dagger}(k),\hat{f}(k)\mapsto\hat{a}^{\dagger}_{k},\hat{a}_{k}$, and $\hat{g}^{\dagger}(k),\hat{g}(k)\mapsto\hat{b}_{k},\hat{b}^{\dagger}_{k}$. Therefore, the general solution to the Dirac equation in Rindler space-time is [20, 41]
$\hat{\psi}(\tilde{t},\rho)=\int^{\infty}_{-\infty}d\mu(k)\left(\hat{a}_{k}e^{-i|k|\tilde{t}}\psi_{k}(\rho)+\hat{b}^{\dagger}_{k}e^{i|k|\tilde{t}}\phi_{k}(\rho)\right),$ (B.36)
where $\hat{a}_{k}=\langle\psi_{k}(\rho),\psi(0,\rho)\rangle$, $\hat{b}^{\dagger}_{k}=\langle\phi_{k}(\rho),\psi(0,\rho)\rangle$, and $\{\hat{a}_{k},\hat{a}^{\dagger}_{q}\}=\{\hat{b}_{k},\hat{b}^{\dagger}_{q}\}=\delta(k-q)$, whilst all other anti-commutation relations vanish.
@@ 4 Quantum field theory of spinor in Minkowski space-time
In Minkowski space-time, as the zweibein is trivially given by $e^{a}_{\mu}=\mathbb{I}$, the Dirac equation, as is given in Eq. (B.14), simplifies to
$\left(i\gamma^{a}e^{\mu}_{a}\partial_{\mu}-m\right)\Psi(T,X)=0$ (B.39)
@@
for the two-component spinor $\Psi(T,X)$. Analogous to the Rindler metric, as the Minkowski metric is stationary with respect to the Minkowski time coordinate $T$, the vector $K^{\mu}=\delta^{\mu}_{T}$ (or equivalently $\partial_{T}$) is a Killing vector that is time-like for all of the Minkowski space-time. Thus, the energy $E$ of the spinor is a well-defined quantity that is conserved along the Minkowski metric’s geodesics [48, 35], and the Dirac equation can be solved via the separation of variables
$\Psi(T,X)=e^{-iE_{K}T}\Psi_{E_{K}}(X),$ (B.40)
@@
where $\Psi_{E_{K}}(X)=(\Psi_{E_{K}}^{+}(X)\ \Psi_{E_{K}}^{-}(X))^{T}$ is a two-component spinor that depends only on $X$. Therefore, the Minkowski space-time Dirac equation can be expressed as the following eigenvalue equation:
$h_{M}\Psi_{E_{K}}(X)=E_{K}\Psi_{E_{K}}(X),$ (B.41)
@@
where $h_{M}=i\gamma^{0}\gamma^{1}\partial_{X}+m\gamma^{0}$ is a linear differential operator that is Hermitian under the conserved inner product
$\langle\Phi(X),\Psi(X)\rangle=\int d\mu(X)\Phi^{\dagger}(X)\Psi(X),$ (B.42)
@@
and $d\mu(X)=dX$ is a measure of integration. That is, $h_{M}$ satisfies $\langle\Phi,h_{M}\Psi\rangle=\langle h_{M}\Phi,\Psi\rangle$ [24]. Once again restricting ourselves to the case of a massless spinor and choosing the chiral representation of the $\gamma$ matrices, the Hermitian operator $h_{M}$ can be expressed as $h_{M}=-u\sigma^{z}\partial_{X}\Psi_{E_{K}}(X)$. Therefore, the eigenvalue equation corresponds to the two following decoupled differential equations:
$E_{K}\Psi_{E_{K}}^{\pm}(X)\pm i\partial_{X}\Psi_{E_{K}}^{\pm}(x)=0,$ (B.43)
@@
where $\Psi_{E_{K}}^{\pm}(x)$ denote the components of the spinor $\Psi_{E_{K}}(x)$ that have the solutions $\Psi_{E_{K}}^{\pm}(X)=\mathcal{N}e^{\pm iE_{K}X}$, where $\mathcal{N}=1/\sqrt{2\pi}$ is a normalisation constant chosen such that $\langle\Psi_{E_{K}}(X),\Psi_{E_{Q}}(X)\rangle=\delta(K-Q)$. Hence, as a solution to the Dirac equation in Minkowski space-time, we have
$\Psi(T,X)=\frac{1}{\sqrt{2\pi}}e^{-iE_{K}T}\begin{pmatrix}e^{iE_{K}X}\\ e^{-iE_{K}X}\end{pmatrix}$ (B.44)
@@
The charge conjugate solution $\Phi(T,X)=C\Psi^{*}(T,X)=e^{iET}\Phi_{E_{K}}(X)$ to the Minkowski space-time Dirac equation, which satisfies the eigenvalue equation $h_{M}\Phi_{E_{K}}(x)=-E_{K}\Phi_{E_{K}}(X)$ and the orthonormality condition $\langle\Phi_{E_{K}},\Psi_{E_{K}}\rangle=0$, will be given by $\Phi(T,X)=\Psi^{*}(T,X)$.
@@
Again, assuming that the orthonormal basis formed by the eigenfunctions $\Psi_{E_{K}}(X)$ of the Hamiltonian $h_{M}$ is complete, any function $F(X)$ within the Hilbert space $\mathcal{H}_{M}$ defined by the conserved inner product of Eq. (B.42) can be decomposed into a continuous sum of eigenfunctions $\Psi_{E_{K}}(X)$ as [23, 24, 25]
$F(X)=\int d\mu(K)\tilde{f}(K)\Psi_{E_{K}}(X),$ (B.45)
@@
$\langle\tilde{f}(K),\tilde{g}(K)\rangle=\int\mu(K)\tilde{f}^{\dagger}(K)\tilde{g}(K),$ (B.46)
@@
of a corresponding Hilbert space $\mathcal{L}^{2}(K)$ of functions $\tilde{f}(K)$, which is related to $\mathcal{H}_{M}$ via the eigenfunction decomposition of Eq. (B.45) and its inverse [23, 24, 25]
$\tilde{f}(K)=\int d\mu(X)\Psi_{E_{K}}^{\dagger}(X)F(X).$ (B.47)
@@
To consistently relate the two Hilbert spaces via the eigenfunction decomposition and its inverse, we require the normalisation conditions $\langle\Psi_{E_{K}}(X),\Psi_{Q}(X)\rangle=\delta(K-Q)$, where $\int d\mu(Q)\delta(K-Q)\tilde{f}(Q)=\tilde{f}(K)$, and $\langle\Psi_{E_{K}}(X),\Psi_{E_{K}}(X^{\prime})\rangle=\delta(X-X^{\prime})$, where $\int d\mu(X^{\prime})\delta(X-X^{\prime})F(X^{\prime})=F(X)$ [23, 24, 25].
@@
As in the Rindler case, from the eigenfunction decomposition of Eq. (B.45), the most general solution $\Psi(T,X)$ of the Minkowski space-time Dirac equation can be decomposed into a continuous sum of the Dirac equation’s spinor $\Psi_{E_{K}}(X)$ and charge conjugate spinor $\Phi_{E_{K}}(X)$ eigenfunctions. That is,
$\begin{split}\Psi(T,X)=\int^{\infty}_{-\infty}d\mu(K)\Bigl(&\tilde{f}(K)e^{-i|K|T}\Psi_{K}(X)\\ &+\tilde{g}(K)e^{i|K|T}\Phi_{K}(X)\Bigr),\end{split}$ (B.48)
@@
where, to once again ensure that the spinor $\Psi_{E_{K}}$ and its charge conjugate $\Phi_{E_{K}}$ always correspond to positive and negative energy solutions, respectively, we have defined
$K=\begin{cases}E_{K},&\text{if }E_{K}\geq 0\\ -E_{K},&\text{if }E_{K}<0\end{cases},\ u_{K}=\begin{cases}u_{+},&\text{if }K\geq 0\\ u_{-},&\text{if}K<0\end{cases}$ (B.49)
and the eigenfunctions $\Psi_{K}(X)=\mathcal{N}e^{-iKX}u_{K}$ and $\Phi_{K}(X)=\mathcal{N}e^{iKX}u_{K}$.
@@
Once again, this solution to the Dirac equation for a spinor in Minkowski space-time can be quantised by promoting $\Psi(T,X)$ and $\Pi(T,X)=-\frac{i}{2}\Psi^{\dagger}(T,X)$ to Hermitian operators and imposing the equal-time canonical anti-commutations of Eq. (A.19). In quantising the Dirac field in Minkowski space-time, the functions $\tilde{f}(K)$ and $\tilde{g}(K)$ are promoted to the operators $\hat{f}(K)$ and $\hat{g}(K)$, respectively. Using the inverse eigenfunction decomposition defined in (B.47), it can be verified that the operators $\hat{f}(K)$ and $\hat{g}(K)$ can be expressed in terms of the Cauchy initial data, $\hat{\Psi}(0,X)$ and $\hat{\Pi}(0,X)$, as $\hat{f}(K)=\langle\hat{\Psi}_{K}(X),\hat{\Psi}(0,X)\rangle$ and $\hat{g}(K)=\langle\hat{\Phi}_{K}(X),\hat{\Psi}(0,X)\rangle$. From these expressions and the orthonormality conditions of $\hat{\Psi}_{K}(X)$ and $\hat{\Phi}_{K}(X)$, the following anti-commutation relations can be derived [48]:
$\{\hat{f}(K),\hat{f}^{\dagger}(Q)\}=\{\hat{g}(K),\hat{g}^{\dagger}(Q)\}=\delta(K-Q),$ (B.50)
$\{\hat{f}(K),\hat{f}(Q)\}=\{\hat{g}(K),\hat{g}^{\dagger}(Q)\}=0,$ (B.51)
@@
Analogous to the Rindler space-time case, it then follows that $\hat{f}^{\dagger}(K)$ and $\hat{g}(K)$ are operators that act to raise and lower the energy of the quantum field by creating and annihilating quanta $\Psi_{K}$ of energy $|K|/2$, respectively, whilst $\hat{g}^{\dagger}(K)$ and $\hat{g}(K)$ are operators that act to lower and raise the energy of the quantum field by creating and annihilating quanta $\Phi_{K}$ of energy |K|/2-|K|/2, respectively. Thus, after re-labelling $\hat{f}(K),\hat{f}(K)\mapsto\hat{A}^{\dagger}(K),\hat{A}_{K}$, and $\hat{g}^{\dagger}(K),\hat{g}(K)\mapsto\hat{B}_{K},\hat{B}^{\dagger}_{K}$, the general solution to the Dirac equation in Minkowski space-time is [41]
$\begin{split}\hat{\Psi}(T,X)=\int^{\infty}_{-\infty}d\mu(K)\Bigl(&\hat{A}_{K}e^{-i|K|T}\Psi_{K}(X)\\ &+\hat{B}^{\dagger}_{K}e^{i|K|T}\Phi_{K}(X)\Bigr),\end{split}$ (B.52)
where $\hat{A}_{K}=\langle\Psi_{K},\Psi(0,X)\rangle$, $\hat{B}^{\dagger}_{K}=\langle\Phi_{K},\Psi(0,X)\rangle$, and $\{\hat{A}_{K},\hat{A}^{\dagger}_{Q}\}=\{\hat{B}_{K},\hat{B}^{\dagger}_{Q}\}=\delta(K-Q)$, whilst all other anti-commutation relations vanish.